OpenStudy (anonymous):

A student club has 11 members. How many ways can the club choose a president, vice president, and treasurer? A. 330 B. 990 C. 165 D. 1,650

OpenStudy (slaaibak):

Solve this by asking yourself: How many people can be chosen for president: After a president have been chosen, how many are left standing for vice president? After you chose the vice, how many remain for the treasurer? Use that logic to solve it

jimthompson5910 (jim_thompson5910):

P(n,r) = (n!)/((n-r)!) P(11,3) = (11!)/((11-3)!) P(11,3) = (11!)/(8!) P(11,3) = (11*10*9*8!)/(8!) P(11,3) = 11*10*9 P(11,3) = 990 So there are 990 ways